Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2025 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 76     Entry time: 30 May 2025, 07:16
Author: Yoonhee Kim 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 9.3 keV, 2000 uJ pulse energy. Beam was unstable at the beginning.

Optical laser delivery

  • ns laser for particle illumination inside the chamber,
  • another ns laser to visualize the tip of capillary on top of the chamber

Achievements/Observations

  • Measured GV for ~ 2 hours

Issues

  • PSI from the first reservoir was leaking
  • Concentration of PSI seems low (judged by seeing camera for liquid jet delivery).
  • We could not see any diffraction from PSI by both AGIPD preview and OnDA (same setting with GV used)

Instrument status

  • AGIPD HV off.
  • AGIPD gate valve closed and pinned.
  • Attenuators set to thickest diamond.
  • Vacuum valves closed.
  • AGIPD breaks enguaged
  • ns Lasers off
ELOG V3.1.4-7c3fd00